# Inverse Purcell Suppression of Decoherence in Majorana Qubits via Environmental Engineering

### v1

@@ Abstract
We propose a novel approach for optimizing topological quantum devices: instead of merely isolating qubits from environmental noise, we engineer the environment to actively suppress decoherence. For a Majorana qubit in a topological superconducting wire, we show that coupling via a parity-flipping operator ($\gamma_{L}$) to a broadband bosonic environment yields a decoherence rate scaling as $\Gamma\propto\rho(\epsilon)/\epsilon$, where $\rho(\epsilon)$ is the environmental density of states at the qubit splitting energy $\epsilon\sim e^{-L/\xi}$. By designing environments with suppressed density of states at low frequencies (e.g., using photonic bandgap materials or Josephson junction arrays), we achieve an “inverse Purcell effect” that leads to a suppression factor $F_{P}\propto(\epsilon/\omega_{c})^{\alpha}$, dramatically reducing decoherence. This provides a quantitative design principle where environmental engineering transforms detrimental noise into a tool for coherence stabilization. Our work establishes environmental engineering as a powerful approach for enhancing topological quantum devices.
@@ I Introduction
The pursuit of fault-tolerant quantum computation has motivated intense research into topological qubits, with Majorana zero modes in superconducting wires emerging as a leading platform Kitaev (2001); Alicea (2012); Sau et al. (2010). Experimental progress in semiconductor-superconductor heterostructures has demonstrated signatures consistent with Majorana zero modes Mourik2012 ; Albrecht2016 , bringing topological quantum computation closer to reality. While topological protection provides inherent resilience against local perturbations, real-world implementations remain susceptible to decoherence from various environmental couplings Schlosshauer (2007), including quasiparticle poisoning Rainis2013 ; Karzig2017 and electromagnetic noise.
@@
Traditional quantum device optimization focuses on maximizing isolation from environmental noise through improved materials and shielding. In this work, we propose and analyze a fundamentally different approach: we demonstrate that the environment itself can be engineered to actively suppress decoherence. This novel approach is motivated by recent advances in quantum circuit engineering that enable unprecedented control over electromagnetic environments Forn-Diaz2019 .
@@
The conventional Purcell effect Purcell (1946) describes how a cavity can enhance emission rates by increasing the local density of states. Here, we demonstrate the opposite phenomenon: by engineering environments with strongly suppressed density of states at the characteristic frequency of topological qubits, we can achieve an “inverse Purcell effect” that actively protects against decoherence. This approach is particularly powerful for Majorana-based qubits, where the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ makes them naturally compatible with frequency-selective environmental engineering using existing platforms like high-impedance resonators Masluk2012 and Josephson junction arrays Pop2014 ; Manucharyan2012 ; Peltonen2013 .
@@
In this Letter, we establish a new design principle for quantum devices: instead of treating environmental coupling as purely detrimental, we show how to harness engineered environments as a resource for coherence protection. By embedding a Majorana qubit in a carefully designed electromagnetic or phononic environment Masluk2012 ; Peltonen2013 ; Safavi-Naeini2014 with suppressed density of states at low frequencies, we transform potential decoherence pathways into tools for coherence stabilization. This approach complements recent theoretical proposals for quasiparticle-poisoning protection Karzig2017 and provides a versatile experimental pathway for enhancing topological quantum devices through environmental design.
@@
While suppression of spontaneous emission through environmental engineering has been demonstrated in various contexts Yablonovitch1987 ; Hulet:1985zz , the application to topological quantum devices and the specific quantitative framework developed here are novel. The unique energy scales of Majorana qubits $\epsilon\sim e^{-L/\xi}$ and their parity-based encoding are specifically tailored to exploit topological protection, which makes them more effective than generic approaches.
@@ II Model of the Majorana Qubit and Environment
We consider a topological superconducting wire of length $L$ hosting Majorana zero modes $\gamma_{L}$ and $\gamma_{R}$ at its ends. The qubit is encoded in the fermion parity basis, where the two degenerate ground states $\{|0\rangle,|1\rangle\}$ are distinguished by their total fermion parity. These states correspond to the occupation of the non-local fermion mode $f=(\gamma_{L}+i\gamma_{R})/2$, such that $|0\rangle$ is the vacuum state ($f^{\dagger}f|0\rangle=0$) and $|1\rangle$ is the occupied state ($f^{\dagger}f|1\rangle=|1\rangle$). The energy splitting $\epsilon$ between these states is exponentially small, $\epsilon\sim\Delta e^{-L/\xi}$, where $\Delta$ is the superconducting gap and $\xi$ is the coherence length; this renders them degenerate for practical purposes. This exponential suppression provides the topological protection against local perturbations that cannot simultaneously affect both widely separated Majorana modes. For our purposes, the key feature is the exponential dependence on $L/\xi$; the specific prefactor $\Delta$ can be absorbed into the overall energy scale and does not affect the functional scaling relationships that are the focus of this work. The system Hamiltonian for the isolated wire is Kitaev (2001):
@@
The decoherence mechanism arises from environmental coupling via a parity-flipping operator such as $\gamma_{L}$. While the parity operator $i\gamma_{L}\gamma_{R}$ distinguishes the parity states ($\langle 0|i\gamma_{L}\gamma_{R}|0\rangle=-1$ and $\langle 1|i\gamma_{L}\gamma_{R}|1\rangle=+1$), it is diagonal in the parity basis and cannot drive transitions between $|0\rangle$ and $|1\rangle$. However, individual Majorana operators like $\gamma_{L}$ have off-diagonal matrix elements $\langle 0|\gamma_{L}|1\rangle=1$ and $\langle 1|\gamma_{L}|0\rangle=1$, enabling them to drive transitions between the qubit states when coupled to a bosonic environment. The interaction Hamiltonian is:
$H_{\text{int}}=g\int_{0}^{L}dx\,\Phi(x)\gamma_{L}\delta(x-x_{0}),$ (2)
where $g$ is the weak coupling strength (with dimensions [M] in natural units), $\Phi$ is the bosonic field, and we approximate the localized nature of the Majorana modes by taking the coupling at a point $x_{0}$ (e.g., the center of the wire for symmetric coupling). This localized coupling approximates the dominant environmental interaction at a junction or defect site.
The parity-flipping operator $\gamma_{L}$, while built from fermionic Majorana operators, acts as a collective degree of freedom that can couple to bosonic environmental modes. This mechanism is analogous to how individual fermions in atoms or quantum dots emit photons through electronic transitions. The bosonic field $\Phi$ represents environmental modes. In the Schrödinger picture (at time $t=0$), its expansion is:
$\Phi(x)=\sum_{k}\frac{1}{\sqrt{2\omega_{k}L}}\left(b_{k}e^{ikx}+b_{k}^{\dagger}e^{-ikx}\right)\,.$ (3)
@@
where $b_{k}^{\dagger}$ and $b_{k}$ are creation and annihilation operators for environmental bosons with frequency $\omega_{k}$. The full Hamiltonian of the combined system is then $H=H_{\text{wire}}+H_{\text{env}}+H_{\text{int}}$, where $H_{\text{env}}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the Hamiltonian of the free environment.
@@
This localized coupling model employs a delta-function approximation $\delta(x-x_{0})$ to represent coupling concentrated at a specific location, such as a junction or defect site. This common theoretical simplification captures the essential physics while enabling analytical treatment. Although more distributed coupling would yield quantitative differences, the fundamental scaling relationship $\Gamma\propto\rho(\epsilon)/\epsilon$ derived here captures the key dependence on environmental density of states.
III Decoherence Rate Calculation
The matrix element for the transition between qubit states with emission of an environmental boson is derived using the full quantized field. For a transition from initial state $|i\rangle=|1\rangle\otimes|0_{\Phi}\rangle$ (excited qubit, vacuum environment) to final state $|f\rangle=|0\rangle\otimes|1_{k}\rangle$ (ground-state qubit, one boson in mode $k$, where $|1_{k}\rangle=b_{k}^{\dagger}|0_{\Phi}\rangle$), the matrix element is obtained from the interaction Hamiltonian (2):
$\langle f|H_{\text{int}}|i\rangle$ $=g\langle 0_{\Phi}|\langle 0|\int_{0}^{L}dx\,\Phi(x)\gamma_{L}\delta(x-x_{0})|1\rangle|0_{\Phi}\rangle$
$=g\langle 0|\gamma_{L}|1\rangle\langle 0_{\Phi}|b_{k}\Phi(x_{0})|0_{\Phi}\rangle.$ (4)
The fermionic matrix element evaluates to:
$\langle 0|\gamma_{L}|1\rangle=1\,$, (5)
since $\gamma_{L}$ acts as a parity-flipping operator that connects the two qubit states. The bosonic matrix element yields:
$\langle 0_{\Phi}|b_{k}\Phi(x_{0})|0_{\Phi}\rangle=\frac{1}{\sqrt{2\omega_{k}L}}e^{-ikx_{0}}\,.$ (6)
Thus, the complete matrix element is:
$\langle f|H_{\text{int}}|i\rangle=\frac{g}{\sqrt{2\omega_{k}L}}e^{-ikx_{0}}\,.$ (7)
The squared magnitude is:
$|\langle f|H_{\text{int}}|i\rangle|^{2}=\frac{g^{2}}{2\omega_{k}L}\,.$ (8)
The matrix element $\langle f|H_{\text{int}}|i\rangle$ is evaluated in the Schrödinger picture and is time-independent. The transition rate $\Gamma$ is given by Fermi’s Golden Rule, which involves the squared magnitude of this matrix element and ensures energy conservation.
Applying Fermi’s Golden Rule, the decoherence rate becomes:
$\Gamma=2\pi\sum_{k}|\langle f|H_{\text{int}}|i\rangle|^{2}\delta(\epsilon-\omega_{k})=\pi\sum_{k}\frac{g^{2}}{\omega_{k}L}\delta(\epsilon-\omega_{k})\,.$ (9)
Converting the sum to an integral over density of states $\rho(\omega)$, we obtain:
$\Gamma=\frac{\pi g^{2}}{L}\int\frac{\rho(\omega)}{\omega}\delta(\epsilon-\omega)\,d\omega=\frac{\pi g^{2}}{L}\frac{\rho(\epsilon)}{\epsilon}\,.$ (10)
For a 1D environment with linear dispersion $\omega_{k}=v|k|$, the free-space density of states is $\rho_{\text{free}}(\omega)=L/(\pi v)$, yielding:
$\Gamma_{\text{free}}=\frac{g^{2}}{v\epsilon}\,.$ (11)
@@ III Decoherence Rate Calculation
This result reveals the fundamental challenge: for topological qubits with exponentially small $\epsilon\sim e^{-L/\xi}$, the decoherence rate diverges as $\Gamma\propto 1/\epsilon$ in a standard environment. Environmental engineering must therefore aim to suppress $\rho(\epsilon)$ rather than enhance it.
This calculation assumes weak coupling where first-order perturbation theory (Fermi’s Golden Rule) is valid, which is appropriate for the parameter regimes considered here.
@@ IV Broadband Purcell Suppression for Coherence Enhancement
The engineered decoherence rate, found by collecting Eq. (10), is:
$\Gamma_{\text{engineered}}=\frac{\pi g^{2}}{L}\frac{\rho_{\text{engineered}}(\epsilon)}{\epsilon}\,.$ (13)
Substituting $\rho_{\text{engineered}}(\epsilon)=\rho_{\text{free}}(\epsilon)(\epsilon/\omega_{c})^{\alpha}$ and $\rho_{\text{free}}(\epsilon)=L/(\pi v)$, we obtain:
$\Gamma_{\text{engineered}}=\frac{g^{2}}{v}\epsilon^{\alpha-1}\omega_{c}^{-\alpha}\,.$ (14)
Since $\epsilon\sim e^{-L/\xi}$, the $L$-dependence scales as:
$\Gamma_{\text{engineered}}\propto\epsilon^{\alpha-1}\propto e^{-(\alpha-1)L/\xi}\,.$ (15)
This scaling reveals a crucial insight: if $\alpha>1$ (strong suppression of density of states at low frequencies), then $\Gamma_{\text{engineered}}$ decreases exponentially with $L$, meaning that the coherence time increases with wire length. This means longer wires can have better coherence protection—a counterintuitive result that aligns with topological protection. If $\alpha=1$, $\Gamma_{\text{engineered}}$ is independent of $L$. If $\alpha<1$, $\Gamma_{\text{engineered}}$ increases with $L$, but this case is less desirable.
This underscores the power of environmental engineering: by choosing materials or structures with large $\alpha$ (e.g., photonic bandgaps with sharp cutoffs), we can make decoherence rates decrease with wire length, enhancing topological protection.
The resulting suppression factor compared to free space, found by collecting Eqs. (11) and (14), is:
$F_{P}=\frac{\Gamma_{\text{engineered}}}{\Gamma_{\text{free}}}=\frac{\rho_{\text{engineered}}(\epsilon)}{\rho_{\text{free}}(\epsilon)}=(\epsilon/\omega_{c})^{\alpha}\,.$ (16)
[…]
$F_{P}\sim(0.1)^{\alpha}\ll 1\,.$ (17)
providing dramatic suppression of decoherence. This “inverse Purcell” effect transforms environmental coupling from a decoherence liability into a coherence-stabilizing asset.

### v2

@@ Abstract
We propose a novel approach for optimizing topological quantum devices: instead of merely isolating qubits from environmental noise, we engineer the environment to actively suppress decoherence. For a Majorana qubit in a topological superconducting wire, the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ provides protection against local perturbations but renders it highly susceptible to pure dephasing from low-frequency environmental noise. We show that coupling via a parity-conserving operator ($i\gamma_{L}\gamma_{R}$) to a bosonic environment yields a dephasing rate $\Gamma_{\phi}\propto S(\epsilon)$, where $S(\epsilon)$ is the environmental noise power at the qubit splitting frequency. In the experimentally relevant regime where $k_{B}T\gtrsim\hbar\epsilon$ (with $T\sim 10-100$ mK), the noise power scales as $S(\epsilon)\propto\rho(\epsilon)k_{B}T/\hbar\epsilon$, leading to a dephasing rate $\Gamma_{\phi}\propto\rho(\epsilon)T/\epsilon$. This exposes a fundamental challenge: the dephasing rate diverges as $1/\epsilon$ for a standard environment, e.g., a 1D system with linear dispersion where $\rho(\epsilon)$ is constant. We overcome this by designing environments with a suppressed density of states following $\rho_{\text{engineered}}(\epsilon)=\rho_{\text{free}}(\epsilon)(\epsilon/\omega_{c})^{\alpha}$. This creates an “inverse Purcell effect” that yields a temperature-independent suppression factor $F_{P}=(\epsilon/\omega_{c})^{\alpha}$. For $\alpha>1$, the engineered dephasing rate decreases exponentially with wire length, $\Gamma_{\phi,\text{engineered}}\propto e^{-(\alpha-1)L/\xi}$, meaning longer wires provide better coherence protection. This provides a quantitative design principle where environmental engineering transforms detrimental noise into a tool for coherence stabilization, while respecting fermion parity superselection rules. Our work establishes environmental engineering as a powerful approach for enhancing topological quantum devices.
@@ I Introduction
The pursuit of fault-tolerant quantum computation has motivated intense research into topological qubits, with Majorana zero modes in superconducting wires emerging as a leading platform Kitaev2001 ; Alicea2012 ; Sau2010 . Experimental progress in semiconductor-superconductor heterostructures has demonstrated signatures consistent with Majorana zero modes Mourik2012 ; Albrecht2016 , bringing topological quantum computation closer to reality. While topological protection provides inherent resilience against local perturbations through exponentially small energy splittings $\epsilon\sim e^{-L/\xi}$, this very property creates a fundamental vulnerability: such small splittings render the qubits highly susceptible to pure dephasing from low-frequency environmental noise. Real-world implementations therefore remain susceptible to decoherence from various environmental couplings Schlosshauer2007 , including quasiparticle poisoning Rainis2013 ; Karzig2017 (a relaxation error) and broadband electromagnetic noise (a primary source of dephasing).
@@
Traditional quantum device optimization focuses on maximizing isolation from environmental noise through improved materials and shielding. In this work, we propose and analyze a fundamentally different approach: we demonstrate that the environment itself can be engineered to actively suppress decoherence by targeting this specific vulnerability. This novel approach is motivated by recent advances in quantum circuit engineering that enable unprecedented control over electromagnetic environments Forn-Diaz2019 .
@@
The conventional Purcell effect Purcell1946 describes how a cavity can enhance emission and relaxation rates by increasing the local density of states. Here, we demonstrate an analogous but opposite phenomenon for dephasing: by engineering environments with strongly suppressed density of states at the specific frequencies relevant to the qubit’s energy splitting, we achieve an “inverse Purcell effect” that actively protects against phase randomization. This approach is particularly powerful for Majorana-based qubits, where the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ makes them naturally compatible with frequency-selective environmental engineering. Furthermore, their parity-based encoding requires coupling via a parity-conserving operator ($i\gamma_{L}\gamma_{R}$), which intrinsically suppresses relaxation and isolates dephasing as the dominant decoherence channel to target. Existing platforms like high-impedance resonators Masluk2012 and Josephson junction arrays Pop2014 ; Manucharyan2012 ; Peltonen2013 provide the necessary tools to implement this spectral shaping.
@@
In this Letter, we establish a new design principle for quantum devices: instead of treating environmental coupling as purely detrimental, we show how to harness engineered environments as a resource for coherence protection. By embedding a Majorana qubit in a carefully designed electromagnetic or phononic environment Masluk2012 ; Peltonen2013 ; Safavi-Naeini2014 with a suppressed density of states around its transition frequency $\epsilon$, we transform the vulnerability to low-frequency noise into a tool for coherence stabilization. Specifically, we demonstrate that proper environmental engineering can yield an exponential suppression of dephasing rates with increasing wire length ($\Gamma_{\phi}\propto e^{-(\alpha-1)L/\xi}$) and create temperature-independent protection. This approach complements recent theoretical proposals for quasiparticle-poisoning protection Karzig2017 and provides a versatile experimental pathway for enhancing topological quantum devices through environmental design.
@@
While suppression of spontaneous emission through environmental engineering has been demonstrated in various contexts Yablonovitch1987 ; Hulet:1985zz , the application to topological quantum devices and the specific quantitative framework we develop for dephasing suppression are novel. The unique energy scales of Majorana qubits $\epsilon\sim e^{-L/\xi}$ and their parity-based encoding are specifically tailored to exploit this form of protection, making the environmental engineering approach more effective for them than for generic qubit systems.
@@ II Model of the Majorana Qubit and Environment
We consider a topological superconducting wire of length $L$ hosting Majorana zero modes $\gamma_{L}$ and $\gamma_{R}$ at its ends. The qubit is encoded in the fermion parity basis, where the two degenerate ground states $\{|0\rangle,|1\rangle\}$ are distinguished by their total fermion parity. These states correspond to the occupation of the non-local fermion mode $f=(\gamma_{L}+i\gamma_{R})/2$, such that $|0\rangle$ is the vacuum state ($f^{\dagger}f|0\rangle=0$) and $|1\rangle$ is the occupied state ($f^{\dagger}f|1\rangle=|1\rangle$). The energy splitting $\epsilon$ between these states is exponentially small, $\epsilon\sim\Delta e^{-L/\xi}$, where $\Delta$ is the superconducting gap and $\xi$ is the coherence length; this renders them degenerate for practical purposes. This exponential suppression provides the topological protection against local perturbations that cannot simultaneously affect both widely separated Majorana modes. However, this very smallness creates a fundamental vulnerability: the qubit operates at microwave frequencies ($\epsilon\sim 0.1-10$ GHz) where environmental noise is significant, making it highly susceptible to dephasing. For our purposes, the key feature is the exponential dependence on $L/\xi$; the specific prefactor $\Delta$ can be absorbed into the overall energy scale and does not affect the functional scaling relationships that are the focus of this work. The system Hamiltonian for the isolated wire is Kitaev2001 :
@@
The decoherence mechanism arises specifically from environmental coupling that causes pure dephasing. We use a parity-conserving operator $i\gamma_{L}\gamma_{R}$ that commutes with the total fermion parity and thus cannot cause transitions between parity sectors; instead, it modulates the energy splitting between states $|0\rangle$ and $|1\rangle$, leading exclusively to dephasing. The interaction Hamiltonian is:
$H_{\text{int}}=g\int_{0}^{L}dx\,\Phi(x)(i\gamma_{L}\gamma_{R})\delta(x-x_{0})\,$, (2)
where $g$ is the weak coupling strength, $\Phi$ is the bosonic field, and we approximate localized coupling at a point $x_{0}$ (e.g., the center of the wire for symmetric coupling). This choice respects the fermion parity superselection rule Wick:1970bd ; Hassler2014 , as $i\gamma_{L}\gamma_{R}$ is a bosonic operator built from an even number of fermionic operators.
The bosonic field $\Phi$ represents environmental modes that will cause dephasing through their low-frequency components. In the Schrödinger picture (at time $t=0$), its expansion is:
$\Phi(x)=\sum_{k}\frac{1}{\sqrt{2\omega_{k}L}}\left(b_{k}e^{ikx}+b_{k}^{\dagger}e^{-ikx}\right),$ (3)
@@
where $b_{k}^{\dagger}$ and $b_{k}$ are creation and annihilation operators for environmental bosons with frequency $\omega_{k}$. For a 1D environment with linear dispersion $\omega_{k}=v|k|$, the density of states per unit frequency is constant, $\rho(\omega)=L/(\pi v)$. This density of states directly determines the noise power spectrum $S(\omega)$ that governs the dephasing rate, as will be shown in the next section. The full Hamiltonian of the combined system is then $H=H_{\text{wire}}+H_{\text{env}}+H_{\text{int}}$, where $H_{\text{env}}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the Hamiltonian of the free environment.
@@
This localized coupling model employs a delta-function approximation $\delta(x-x_{0})$ to represent coupling concentrated at a specific location, such as a junction or defect site. This common theoretical simplification captures the essential physics while enabling analytical treatment. Although more distributed coupling would yield quantitative differences, the fundamental scaling relationship $\Gamma_{\phi}\propto S(\epsilon)$ derived here captures the key dependence on environmental noise power, particularly the critical role of low-frequency noise due to the small energy splitting $\epsilon$.
III Dephasing Rate Calculation
The dephasing rate for a qubit with energy splitting $\epsilon$ is determined by the noise power spectrum $S(\omega)$ of the environment. For the parity-conserving interaction $H_{\text{int}}=g(i\gamma_{L}\gamma_{R})\Phi$, which commutes with the system Hamiltonian, the dephasing rate follows the standard result for pure dephasing Schoelkopf2003 ; Clerk2010 :
$\Gamma_{\phi}=\frac{g^{2}}{2}S(\epsilon),$ (4)
where $S(\epsilon)$ is the noise power at frequency $\epsilon$. This expression derives from the mode expansion of $\Phi$ and the fluctuation-dissipation theorem Callen1951 ; Kubo1966 , which relates $S(\omega)$ to the density of states $\rho(\omega)$ and temperature. For a bath of harmonic oscillators, this yields Clerk2010 ; Gardiner2004 :
$S(\omega)=2\pi\rho(\omega)\left[n(\omega)+\frac{1}{2}\right]=\pi\rho(\omega)\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),$ (5)
where $n(\omega)=(e^{\hbar\omega/k_{B}T}-1)^{-1}$ is the Bose-Einstein distribution. In the low temperature limit ( $k_{B}T\ll\hbar\omega$), the noise power simplifies to $S(\omega)\approx\pi\rho(\omega)$, while in the high-temperature limit ($k_{B}T\gg\hbar\omega$), it scales as $S(\omega)\approx 2\pi\rho(\omega)\frac{k_{B}T}{\hbar\omega}$.
For Majorana qubits with $\epsilon\sim 1$ GHz ($\hbar\epsilon/k_{B}\sim 48$ mK) and operating at temperatures $T\sim 10-100$ mK, we typically satisfy $k_{B}T\gtrsim\hbar\epsilon$. This places us in the regime where the high-temperature approximation applies:
$S(\epsilon)\approx 2\pi\rho(\epsilon)\frac{k_{B}T}{\hbar\epsilon}\,$, (6)
and consequently, the dephasing rate becomes:
$\Gamma_{\phi,\text{free}}\approx\frac{\pi g^{2}k_{B}T}{\hbar\epsilon}\rho_{\text{free}}(\epsilon)\,.$ (7)
@@ III Dephasing Rate Calculation
This result reveals the fundamental challenge: for topological qubits with exponentially small $\epsilon\sim e^{-L/\xi}$, the dephasing rate in a standard environment diverges as $\Gamma_{\phi,\text{free}}\propto 1/\epsilon$ because the thermal factor $k_{B}T/(\hbar\epsilon)$ grows as $1/\epsilon$ while $\rho_{\text{free}}(\epsilon)$ remains approximately constant for a 1D environment with linear dispersion $\omega_{k}=v|k|$. Environmental engineering must therefore aim to suppress $\rho(\epsilon)$ rather than enhance it to counteract this divergence.
@@ IV Broadband Purcell Suppression for Coherence Enhancement
$\frac{\rho_{\text{engineered}}(\omega)}{\rho_{\text{free}}(\omega)}=\begin{cases}(\omega/\omega_{c})^{\alpha},&\omega<\omega_{c},\\ 1,&\omega\geq\omega_{c},\end{cases}$ (8)
The engineered dephasing rate is determined by the noise power spectrum at the qubit frequency $\epsilon$:
$\Gamma_{\phi,\text{engineered}}=\frac{g^{2}}{2}S(\epsilon).$ (9)
For a bosonic environment, the noise power spectrum follows from the fluctuation-dissipation theorem:
$S(\epsilon)=\pi\rho(\epsilon)\coth\left(\frac{\hbar\epsilon}{2k_{B}T}\right).$ (10)
In the experimentally relevant regime where $k_{B}T\gtrsim\hbar\epsilon$ (typical for Majorana qubits with $\epsilon\sim 1$ GHz and $T\sim 10-100$ mK), we have:
$S(\epsilon)\approx 2\pi\rho(\epsilon)\frac{k_{B}T}{\hbar\epsilon},$ (11)
leading to the engineered dephasing rate:
$\Gamma_{\phi,\text{engineered}}\approx\frac{\pi g^{2}k_{B}T}{\hbar\epsilon}\rho_{\text{engineered}}(\epsilon).$ (12)
Using the engineered spectral density, $\rho_{\text{engineered}}(\epsilon)=\rho_{\text{free}}(\epsilon)(\epsilon/\omega_{c})^{\alpha}$, and since $\rho_{\text{free}}(\epsilon)=L/(\pi v)$ is approximately constant for small $\epsilon$, we obtain:
$\Gamma_{\phi,\text{engineered}}\propto\frac{g^{2}k_{B}T}{\hbar}\epsilon^{\alpha-1}\omega_{c}^{-\alpha}.$ (13)
Since $\epsilon\sim e^{-L/\xi}$, the $L$-dependence scales as:
$\Gamma_{\phi,\text{engineered}}\propto\epsilon^{\alpha-1}\propto e^{-(\alpha-1)L/\xi}.$ (14)
This scaling reveals a crucial insight: if $\alpha>1$ (strong suppression of density of states at low frequencies), then $\Gamma_{\phi,\text{engineered}}$ decreases exponentially with $L$, meaning that the coherence time increases with wire length. This counterintuitive result—that longer wires provide better coherence protection—aligns perfectly with topological protection.
The resulting suppression factor compared to free space is:
$F_{P}=\frac{\Gamma_{\phi,\text{engineered}}}{\Gamma_{\phi,\text{free}}}=\frac{\rho_{\text{engineered}}(\epsilon)}{\rho_{\text{free}}(\epsilon)}=(\epsilon/\omega_{c})^{\alpha}.$ (15)
Remarkably, the temperature dependence cancels in this ratio, making the suppression factor $F_{P}$ temperature-independent. This is a key advantage of our approach: environmental engineering provides protection against dephasing that is robust against thermal fluctuations in the experimentally relevant regime.
@@
$F_{P}\sim(0.1)^{\alpha}\ll 1,$ (16)
providing dramatic suppression of dephasing. This “inverse Purcell” effect transforms environmental coupling from a decoherence liability into a coherence-stabilizing asset.